use Stokes theorem evaluate ∮A.dr, where A=y²i+x²j+z²k and C is boundary of the part of the plane x+y+z=1 like in the first octant

1
Expert's answer
2021-12-22T03:24:19-0500
"curl \bold A=\begin{vmatrix}
 \bold i & \bold j & \bold k \\
\\
 \dfrac{\partial}{\partial x} & \dfrac{\partial}{\partial y} & \dfrac{\partial}{\partial z} \\
\\
y^2 & x^2 & z^2
\end{vmatrix}"

"=(2x-2y)\bold k"

"\bold n=\langle1, 1, 1\rangle"

"curl \bold A\cdot \bold n=2x-2y"

"\int_C\bold A\cdot d\bold r=\int\int _Scurl \bold A\cdot d\bold S"

"=\int \int_S(2x-2y)dS"

"=\displaystyle\int_{0}^1\displaystyle\int_{0}^{1-y}(2x-2y)dxdy"

"=\displaystyle\int_{0}^1[x^2-2xy]\begin{matrix}
 1-y \\
 0
\end{matrix}dy"

"=\displaystyle\int_{0}^1((1-y)^2-2y(1-y))dy"

"=\displaystyle\int_{0}^1(1-2y+y^2-2y+2y^2)dy"

"=[y^3-2y^2+y]\begin{matrix}
 1 \\
 0
\end{matrix}=0"

"\int_C\bold A\cdot d\bold r=0"